OpenStudy (anonymous):

xy=1 x^2+y^2 = 2 finf the values of x and y

OpenStudy (anonymous):

it's a system of equations, solve the first one for one variable, then substitute it in to the second equation.

hartnn (hartnn):

hi @cassie93 :) there are multiple ways to do this, one of which can be using \((x-y)^2 =... ? \\ (x+y)^2=...?\) and get the values of x-y and x+y , then getting x,y will be very easy

hartnn (hartnn):

or you could also start with what peter said, suit yourself :)

OpenStudy (anonymous):

thnx....i get the idea...!! (x-y)(x+) ....thnx so much...:)

hartnn (hartnn):

good! also tell us what values of x, y you get , so that we can verify with our answers :)

hartnn (hartnn):

because there are more than one pair of x,y values in the solution :)

OpenStudy (anonymous):

i dnt knw hw to solve with 1/y^2....

hartnn (hartnn):

you mean you got 1/y^2 + y^2 = 2 and you're stuck ?

OpenStudy (anonymous):

yup...

OpenStudy (anonymous):

find a common denominator

OpenStudy (anonymous):

then multiply both sides by whatever that common denominator is

hartnn (hartnn):

or if you want to avoid the 4th degree solving, you can assume y^2 = a 1/a +a = 2 multiply 'a' throughout to get a easy quadratic :)

hartnn (hartnn):

when you find values of 'a' , you can resubstitute back a = y^2 to get values of y :)

hartnn (hartnn):

let us know if you have difficulty finding 'a' values

OpenStudy (anonymous):

i dnt undrstand about the y^2 = a 1/a +a = 2

hartnn (hartnn):

we had 1/y^2 + y^2 = 2 right ? thats an equation in y^2 instead of working with y^2, i can take a new variable (here i took 'a') to simplify solving it i just replaced every y^2 by 'a' this is called substitution hence i get 1/a + a = 2

OpenStudy (anonymous):

^what he said

OpenStudy (anonymous):

so, there is two values of 'a'..? i got a=0.4142 and a=-2.4142

OpenStudy (anonymous):

is tht correct?

hartnn (hartnn):

how did you get those ? did you do this ? 1/a + a = 2 1 + a^2 = 2a a^2 -2a +1 =0 ??

OpenStudy (anonymous):

oh, i got a=1

hartnn (hartnn):

that is indeed correct! :) y^2 = 1 then values of y are .. ?

OpenStudy (anonymous):

so, substitute y^2=1 into 1/y^2+y^2=2....?

hartnn (hartnn):

you can actually find the values of y from y^2 =1

hartnn (hartnn):

hint : *cough* take square root on both sides *cough*

OpenStudy (anonymous):

remember that a negative times a negative equals a positive...

OpenStudy (anonymous):

*gives hartnn water*

OpenStudy (anonymous):

so, the values of x=1,-1 y=1,-1 ....??

hartnn (hartnn):

yes!

OpenStudy (anonymous):

exactly right

hartnn (hartnn):

when y=1, you get x=1, right ? so actually write those as (1,1) and (-1,-1)

OpenStudy (anonymous):

TTHHHHAAAAANNNNKKKK UUUUUU PEOPLE......!!! :) U GUYS AWESOME..!!

hartnn (hartnn):

you're awesome too! :D welcome ^_^

OpenStudy (anonymous):

you're the awesome one, we just guided you to the answer

OpenStudy (anonymous):

any other questions?

OpenStudy (anonymous):

thts it so far..:) thnx by the way...:)

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

so where are you from? (the messages aren't working right now)

OpenStudy (anonymous):

Philippine :)

OpenStudy (anonymous):

do you live there now?

OpenStudy (anonymous):

yup...y?

OpenStudy (anonymous):

I am an American living in Beijing, China

OpenStudy (anonymous):

well yes, sort of. my dad works here. I am a senior in high school, so I'm not in college yet.

OpenStudy (anonymous):

you are in college?

OpenStudy (anonymous):

yup, 1st year.. all about refrshing the high school subjects.. and my weakness is in math..

OpenStudy (anonymous):

haha, my strength is math, so I can help you

OpenStudy (anonymous):

cool... now i knw where to find help...:)

OpenStudy (anonymous):

you seem like an interesting person

OpenStudy (anonymous):

#replied u in inbox :)

OpenStudy (anonymous):

Thank you XD (inbox doesn't work very well right now, you have to refresh the page to see if messages have been sent.)

OpenStudy (anonymous):

LOL...ok..just talk here then...:)

OpenStudy (anonymous):

ok, what's your favorite subject?

OpenStudy (anonymous):

Bio... (i'm taking Biological Engineering) :)

OpenStudy (anonymous):

oh wow, you're smart :) biological engineering on the small scale like DNA modification or other stuff like the use of biology in industry?

OpenStudy (anonymous):

yup..exactly...(but not tht smart)... ^_^

OpenStudy (anonymous):

I'm sure you're just being humble.

OpenStudy (anonymous):

my nature. ^_^ wht about u? any plan for college?

OpenStudy (anonymous):

I plan to go back to the US. I plan on studying engineering, either mechanical or petroleum.

OpenStudy (anonymous):

cool...tht suits u... u'r smart, u can take any course u wnt..:) maybe u'll end up in MIT?

OpenStudy (anonymous):

haha no, my grades aren't nearly good enough for that. I'm smart but lazy.

OpenStudy (anonymous):

haha... u'll do fine...

OpenStudy (anonymous):

and on top of that, i'm white. I know only two people who got into MIT, and they were both hispanic (so they had an advantage with admissions) and had better grades than I do.

OpenStudy (anonymous):

hmm... anyway, u'll succeed someday.. :) it's a guarantee.. just embrace ur gift wisely... ^_^

OpenStudy (anonymous):

wnt to ask smthing, is X^3 = 1, then X= +-1 ?

OpenStudy (anonymous):

if x^3 = 1, x is 1, (sqrt(3)/2 + 0.5i) or (sqrt(3)/2 - 0.5i)

OpenStudy (anonymous):

ok...^_^thnx... i got to go... i got Bio quiz 2morrow, need to study... ^_^ nice to meet u Peter... stay connected... ^_^

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

I sent you a message

hartnn (hartnn):

wow you 2 had pretty good convo. :) thank god notification system also isn't working :P else i would've got notifications like every minute :O

OpenStudy (anonymous):

if the notification system was working, the message system would be working too and we would have this convo there

OpenStudy (anonymous):

oh, oops, I got the cube root of one wrong. it's actually 1, -0.5+0.5sqrt(3)i, and -0.5-0.5sqrt(3)i